During rush-hour the equilibrium traffic is 2000 cars. The average private cost of the trip is $5 at the equilibrium traffic level. The marginal social cost of the trip is $9 at the equilibrium traffic level. The efficient traffic level is 1500 cars. The average private cost of the trip is $4 at the efficient traffic level. The marginal social cost of the trip is $7 at the efficient traffic level.

 

  Which traffic volume has a higher average trip speed? 

 

    What is the externality cost per car at the equilibrium traffic level? What is the externality cost per car at the efficient traffic level?

 

    What is the level of the deadweight loss on this highway when it is in equilibrium?

 

    What is the efficient congestion toll on this highway?

 

    If the average fixed cost for expanding the highway is $2, should the state expand the highway? 

 

 

 

1
Expert's answer
2021-03-14T19:52:33-0400

The traffic volume of 1500 cars has a higher average trip speed.

The externality cost per car at the equilibrium traffic level is 9 - 4 = $5. The externality cost per car at the efficient traffic level is 7 - 4 = $3.

The level of the deadweight loss on this highway is 3×1500 = $4500 when it is in equilibrium.

The efficient congestion toll on this highway is $2.

If the average fixed cost for expanding the highway is $2, then the state shouldn't expand the highway.